OpenStudy (anonymous):

Find the Area enclosed by the curve "y=sqrt(5x+4)", x-axis, y-axis, and the line x=1

OpenStudy (amistre64):

the area is integration, a picture is nice but immaterial at the moment

OpenStudy (anonymous):



OpenStudy (amistre64):

 \[A=\int_{a}^{b}f(x)~dx\]

OpenStudy (anonymous):

I got as far as \[\int\limits_{0}^{1} \left( \sqrt{5x+4} \right)dx\]

OpenStudy (amistre64):

convert the rt to a power

OpenStudy (amistre64):

and, if you cant see it clearly, try a u substitution

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\left( 5x+4 \right)^{\frac{ 1 }{ 2}}dx\]

OpenStudy (amistre64):

if we let 5x+4 = u what does dx become?

OpenStudy (anonymous):

i really have no idea

OpenStudy (amistre64):

take a derivative with respect to x ... ive got real life issues pressing so ill be away for a few hours at best